A bushed pin type flexible coupling is used to connect two shafts and transmit 15 kW power at 600 rpm Shafts, keys and pins are made of commercial steel (σyt = σyc = 240 N/mm2) and the factor of safety is 3. The flanges are made of grey cast iron FG 200 (σut = 200 N/mm2) and the factor of safety is 6. Assume, τy = 0.5σyt and τu = 0.5σut. The permissible shear stress for the pins is 35 N/mm2. The permissible bearing pressure for the rubber bushes is 1 N/mm2. The keys have a square cross-section. Design the coupling using manufacturer’s catalogue.
1
Expert's answer
2021-10-26T07:00:51-0400

Permissible stresses;

Shaft, "\tau=\frac{S_{Sy}}{FS}"

"=\frac{0.55*240}{3}=44N/mm^2"

This is same as permissible stress in key

Pins, "\tau=35N/mm^2"

"\sigma_t=\frac{240}{3}=80N/mm^2"

Flanges, "\tau=\frac{S_{Sy}}{FS}"

= "\frac{0.5*200}{6}=16.67N/mm^2"

Diameter of shaft

"P=\frac{2πnM_t}{60}"

"5*10^3=\frac{2π*720M_t}{60}"

"M_t=66314.6Nmm"

"\tau=\frac{16M_t}{πd^3}"

"d^3=\frac{16*66314.6}{π*44}"

d=20mm

Diameter of hub "d_h=2d=40mm"

Length of hub "l_h=1.5d=30mm"

Pitch circle diameter "D=4d=80mm"

Thickness of flange "t=0.5d=10mm"

Thickness of protective rim "t_1=0.25d=5mm"